OpenStudy (lasttccasey):

I need help with this one, Find the derivative by the limit process. f(x) = √(x+4). I have the answer but I can't seem to get it using the difference quotient.

OpenStudy (lasttccasey):

\[\frac{dy}{dx}=\frac{f(x-\Delta x)-f(x)}{\Delta x}\]is the formula I am using. The answer is \[\frac{1}{2\sqrt{x+4}}\]We haven't learned the chain rule yet, its in the next chapter.

OpenStudy (anonymous):

\[ \lim_{h\to0}\frac{\sqrt{x+4+h}-\sqrt{x+4}}{h} \]Multiply by the conjugate? That is to say, \[ \frac{\sqrt{x+4+h}+\sqrt{x+4}}{\sqrt{x+4+h}+\sqrt{x+4}} \]

OpenStudy (anonymous):

That's just my first guess, haven't worked it out yet.

OpenStudy (lasttccasey):

Nvm, I found what I did wrong but yea multiplying by the conjugate is probably easier than what I was going. I tried to square both the top and bottom but I didn't do the top right. Thanks